OpenStudy (anonymous):

8/6 - x, find: x - 7 x

OpenStudy (anonymous):

You're just plugging and chugging here, so plug what ever they have in the brackets as x in your equation.

OpenStudy (anonymous):

So if you had f(1) All you would have to do is this \[f(1) = \frac{ 1 }{ (1)^2 }-3\] or if the 3 is below the denominator what ever it is, you should get the point.

OpenStudy (anonymous):

Anyways, I have to go right now, I'll come back later and check on your work. But it will be very late, so don't wait up :P.

OpenStudy (anonymous):

No, you have to plug in the 3 into where x is.

OpenStudy (anonymous):

That's the point of saying f(3)

OpenStudy (anonymous):

Yes :)

OpenStudy (anonymous):

Yes, solve! Alright, I have to go now, c ya, and good luck!

OpenStudy (jhannybean):

And it'll work the same way for \[f(2+h)=\frac{1}{(2+h)^2}-3\]

OpenStudy (jhannybean):

For this one you just have to expand it :)

OpenStudy (jhannybean):

This one is saying, "if you had a number, let's say 2, and you added a value of h, you'l end up with a function of the increased value"

OpenStudy (jhannybean):

Think of \((h+3)^2\) as a number, since we have a -3 in our function as well, we want to put this over a common denominator, so we multiplythe top and bottom by \( (h+3)^2\)\[\frac{1}{(h+2)^2} - \frac{3}{1} \cdot \frac{(h+2)^2}{(h+2)^2}\]

OpenStudy (jhannybean):

That means we can reduce this to \[\frac{1-3(h+2)^2}{(h+2)^2}\] You see how we did that? :o

OpenStudy (anonymous):

ohhhhh ok so that is all we can do with that

OpenStudy (jhannybean):

No no, we can also expand this function to reduce it even further :P

OpenStudy (jhannybean):

\[ (\color{red}h+\color{blue}2)^2 \implies (\color{red}a+\color{blue}b)^2 = \color{red}a^2+2\color{red}a\color{blue}b +\color{blue}b^2\]

OpenStudy (jhannybean):

In expanding \((h+2)^2\) , we would expand it in that manner, just match up the colors to the format :)

OpenStudy (anonymous):

wait I'mconfused now...

OpenStudy (jhannybean):

a and b is the format of this equation, so a = h and b = 2

OpenStudy (jhannybean):

Since we now know a = h and b= 2, we can use the expanded version of this equation , which is \( \color{red}a^2 + 2\color{red}a\color{blue}b+\color{blue}b^2 \implies \color{red}h^2 +2\cdot \color{red}h\cdot\color{blue}2 +\color{blue}2^2\) Do you see how it's working out? :)

OpenStudy (perl):

f(2+h) = 1/(2+h)^2 -3 agreed?

OpenStudy (perl):